OpenStudy (andijo76):

heights of men on a basketball team have a bell shaped distribution with a mean of 173 cm and a standard deviation of 8 cm using the empirical rule what is the approximate percentage of the men between the following a. 165 cm and 181 cm b. 149 cm and 197cm

OpenStudy (anonymous):

wat exactly is empirical rule?

OpenStudy (mathmale):

andijo: Can you explain to TheCool what the E. Rule signifies? If you could do that, I'll then know better how to respond to your question.

OpenStudy (anonymous):

A I think

OpenStudy (mathmale):

The Empirical Rule (which applies only when your data is reasonably "normal"), allows us to predict what percentage of your data is within 1, 2 and 3 standard deviations from the mean. Andijo? Can you provide those percentages? Within 1 standard dev. of the mean lie approx ?? % of your data. etc.

OpenStudy (mathmale):

Anyone care to share more info about the Empirical Rule?

OpenStudy (andijo76):

i can show you an example i found on yahoo

OpenStudy (anonymous):

If a data distribution is approximately normal then about 68 percent of the data values are within one standard deviation of the mean (mathematically, μ ± σ, where μ is the arithmetic mean), about 95 percent are within two standard deviations (μ ± 2σ), and about 99.7 percent lie within three standard deviations (μ ± 3σ). This is known as the 68-95-99.7 rule, or the empirical rule.

OpenStudy (andijo76):

The empirical rule is also known as the 68 - 95 - 99.7 rule. For normaly distributed data, 68% of your population should fall within 1 standard deviation of the mean. 95% should fall within 2 standard deviations of the mean. 99.7% should fall within 3 dtandard deviations of the mean. In your example, the mean is 168cm. We move one standard deviation in either direction by subtracting and adding 7cm to 168cm. 168-7 = 161. 168+7 = 175 68% of the men on the basketball team have a height between 161cm and 175 cm. Add/subtract another deviation and you get the range for 95% of all values. 161-7 = 154 175+7 = 182 95% of the men on the basketball team have heights between 154 cm and 182 cm.

OpenStudy (mathmale):

Glad you looked this up on the 'Net. Great strategy. Let me give you a hint of what I'm asking you to share: "Within 1 standard dev. of the mean lie approx. 68% of your data." Bet that's what you're seeing on Yahoo.

OpenStudy (mathmale):

Within 2 standard deviations, what percentage of the data lie? Within 3 standard deviations, ??

OpenStudy (mathmale):

Now: You're given two intervals. the first one is: a. 165 cm and 181 cm Given that the mean of this data set is 173, add the std. dev. (8) to 173 to get the upper limit on "1 std. dev. above the mean," and subtract the std. dev. (8) to get the lower limit on "1 std. dev. below the mean."

OpenStudy (mathmale):

Please try doing this. Type in something, even if you're not sure it's "right." Then I could give you more appropriate feedback.

OpenStudy (andijo76):

173-8=165 173+8=181 =68% 165-8=157 175+8=183 =95%

OpenStudy (andijo76):

i think i figured it out with help thanks z1=(165-173)/-1 z2=(181-173)/+1 68% b z1=(149-173)/8=-3 z2=(197-173)/8=+3 95%

OpenStudy (mathmale):

I'm very glad for you! Thanks for your persistence.